OpenStudy (gabylovesyou):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was 3 times his age. How old is Sam now?

OpenStudy (anonymous):

c = 2s c - 10 = 3(s-10) 2s - 10 = 3s - 30 3s - 30 = 2s - 10 3s - 30 - 2s = 2s - 10 -2s s - 30 = -10 s = -10 + 30 s = 20

OpenStudy (anonymous):

Check 20 - 10 = 10, and 40 -10 = 30 10*3 = 30 20 * 2 = 40

OpenStudy (gabylovesyou):

so the answer is 20??

OpenStudy (anonymous):

Yes, Sam is 20 years old now.

OpenStudy (gabylovesyou):

thank u! :)